OpenStudy (anonymous):

Need help solving this small problem. For some reason i'm having a brain fart.

OpenStudy (anonymous):

drawing equation sec

OpenStudy (anonymous):

\[\frac{ X^4-81 }{ x-3 }\]

OpenStudy (anonymous):

factor as the difference of two squares twice

OpenStudy (anonymous):

\[x^4-81=(x^2+9)(x^2-9)=(x^2+9)(x+3)(x-3)\] then cancel

OpenStudy (anonymous):

Alright that makes this easier ty